OpenStudy (anonymous):

is there a technique on solving word problem in age of a person, time consumed in x activity, or any other? im having a hard time solving them during exam while time is so limited

OpenStudy (arindameducationusc):

is it some kind of linear equation problem?

OpenStudy (anonymous):

example a man is 3x as old as his son.4years ago he was 4x as old as his son. how old is his son?

OpenStudy (unklerhaukus):

you could start by changing some of the words to letters [a man is 3x as old as his son] becomes m = 3s

OpenStudy (arindameducationusc):

3x+4=4x

OpenStudy (arindameducationusc):

see, link the current with the old.

OpenStudy (anonymous):

is it really 3x+4=4x? if thats the equation so x=4 and if son is 4years old then 4years ago he was 0

OpenStudy (anonymous):

i guess the right equation is 3x-4 = 4(x-4) while x is the age of son. man is 36 4years ago 32 son is 12 4years ago 8

OpenStudy (arindameducationusc):

if x=4 then its the son's age. Man's age is 3x, which is 3*4=12 But let me check... wait....

OpenStudy (arindameducationusc):

ya this should be the answer.....

OpenStudy (arindameducationusc):

In your equation how did you put4(x-4) ? Just tell me why

OpenStudy (anonymous):

3x-4 is the age of man 4years ago 4(x-4) 4years ago the age of son is 4x as the man

OpenStudy (anonymous):

-4 since its past

OpenStudy (arindameducationusc):

O ho.... see....3x is the man's age. 4x is the age 4 years before. 3x-4 means the age minus 4 years which gives the previous age which is 4x so, 3x-4=4x.. when you find x, you get the son's age... Got it @darkmaker ?

OpenStudy (anonymous):

3x-4=4x -4=4x-3x -4=x since there is no negative age x is equal to 4 if x=4 son is 4 4years ago 0 man is 12 4years ago 8 equation seems wrong age of man is 4x as age of son 4years ago the age of son must be reduced by 4 also since it was 4years ago anyway I browse the last page of the book and there is the correct answer there son is 12 years old =)

OpenStudy (arindameducationusc):

okay good for you then... hmmm let me think where I made the mistake. Thanks for the answer anyways.......

OpenStudy (arindameducationusc):

I got it... You are right......... Good job! @darkmaker

OpenStudy (anonymous):

ill guess im gonna ask my professor regarding my question here. thanks anyway =)